OpenStudy (zenmo):

Find the centroid of the region in the first quadrant bounded by the given curves.

OpenStudy (zenmo):

\[y = x^3, x = y^3\]

OpenStudy (phi):

any thoughts ?

OpenStudy (zenmo):

Find the area in order to set up the integration from a to b, but no idea on how to find them algebraically by setting them equal to each other.

OpenStudy (dominusscholae):

If you the points of intersection, it will give you your bounds. Try substituting the second equation so you can find the y-intersection and then you can rule out answers given that it's in the first quadrant

OpenStudy (phi):

did you sketch the region ?

OpenStudy (zenmo):

No, my graphing skills are very bad, if I were to sketch this -- start sketching this by plotting points?

OpenStudy (phi):

yes, just a few. for y= x^3 try x=0, ½, 1

OpenStudy (zenmo):